# The Automated but Risky Game: Modeling Agent-to-Agent Negotiations and Transactions in Consumer Markets

### v3

@@ Abstract
AI agents are increasingly used in consumer-facing applications to assist with tasks such as product search, negotiation, and transaction execution. In this paper, we investigate a future setting where both consumers and merchants authorize AI agents to automate the negotiations and transactions in consumer settings. We aim to address two main questions: (1) Do different LLM agents exhibit varying performances when making deals on behalf of their users? (2) What are the potential risks when we use AI agents to fully automate negotiations and deal-making in consumer settings? We design an experimental framework to evaluate AI agents’ capabilities and performance in real-world negotiation and transaction scenarios, and experimented with a range of LLM agents. Our analysis reveals that deal-making with LLM agents in consumer settings is an inherently imbalanced game: different AI agents have large disparities in obtaining the best deals for their users. Furthermore, we found that LLMs’ behavioral anomaly might lead to financial loss for both consumers and merchants when deployed in real-world decision-making scenarios, such as overspending or making unreasonable deals. Our findings highlight that while automation can enhance transactional efficiency, it also poses nontrivial risks to consumer markets. Users should be careful when delegating business decisions to LLM agents.
@@ 1 Introduction
Business negotiation and deal-making lie at the heart of the modern economy, yet achieving agreement is rarely straightforward. It requires effective information gathering, strategic reasoning, and, above all, skilled negotiation and decision-making (lewicki2011mastering; agndal2017two). Recently, large language model (LLM) powered AI agents have demonstrated remarkable capabilities and are increasingly being adopted for real-world tasks (xu2024theagentcompany; masterman2024landscape). Given the importance of negotiation and deal-making in business operations, researchers and practitioners have begun exploring ways to leverage AI agents to automate shopping and sales processes for both consumers and merchants (kong2025fishbargain; chen2024chatshop), mostly with an assumption that the agents are interacting with real human users. However, with the rapid adoption of AI agents in consumer markets, both consumers and merchants might delegate their negotiation and decision-making to AI agents and direct agent-to-agent interactions may soon be commonplace. Given the natural capability differences of AI agents in negotiation settings (bianchi2024well) and unique agent-to-agent negotiation dynamics (vaccaro2025advancing), it becomes important to understand what will happen when consumers and merchants use AI Agents with different capabilities to automate their negotiation and transactions in consumer settings.
@@
In this study, we propose a comprehensive framework to investigate the opportunities and risks associated with fully automated, user-authorized agent-to-agent negotiation and transaction. Inspired by real-world shopping and sales workflows, we design an experimental setting in which a buyer agent attempts to negotiate a lower price for a product based on a user-defined budget, while a seller agent, aware of the wholesale cost, aims to maximize its profit. Each agent independently makes its own decisions throughout the negotiation, simulating a fully autonomous, end-to-end transaction between AI agents. To evaluate the negotiation behaviors and capabilities of AI agents in realistic consumer scenarios, we compile a dataset of 100 real-world products across three major categories: electronic devices, motor vehicles, and real estate, covering a diverse set of consumer scenarios with different price ranges. For each item, we collected the actual retail prices and estimated wholesale values, which were then provided to the seller agents to simulate real market dynamics. We conducted negotiation and transaction experiments using several advanced language models, including GPT series (hurst2024gpt), Qwen-2.5 series (yang2024qwen2), and DeepSeek series (liu2024deepseek; guo2025deepseek). Our analysis reveals substantial performance gaps across models. More capable agents consistently outperform their counterparts – earning higher profits when acting as sellers and achieving greater savings as buyers. These disparities emerge not only across different model families but also among models with varying parameter sizes. This suggests that agent-to-agent trading inherently becomes an imbalanced game when agents possess unequal capabilities.
@@
Beyond performance differences, we identify several key risks associated with delegating negotiation and transactional authority to AI agents: (1) Constraint violation risk: Buyer agents may disregard user-imposed budget constraints, completing purchases that the user cannot afford. Similarly, seller agents may accept prices below wholesale costs, leading to sizeable financial losses; (2) Excessive payment risk: buyer agents sometimes offer higher prices than retail price, resulting in unnecessary overpayment; (3) Negotiation deadlock risk: agents may become stuck in prolonged negotiation loops without reaching an agreement; (4) early settlement risk: in high budget settings, the buyer agents tend to compromise more readily despite they are instructed to obtain the lowest possible price. On the contrary, buyer agents with a lower budget are able to obtain significanlty more discounts for the same sets of products.
@@
Our findings have important implications for automated decision-making with AI agents. Access to more powerful AI models can lead to better deals, potentially reinforcing economic disparities among users. Furthermore, weaknesses in LLMs—such as limited numerical reasoning and occasional failures in instruction-following—can expose both consumers and businesses to systemic financial risks. As fully autonomous agent-to-agent interactions become more common, users should be cautious when delegating high-stakes decisions to AI agents. This paper makes the following contributions: (1) We propose a novel and realistic setting for agent-to-agent negotiation and transaction, with clear practical implications for future consumer markets; (2) We design a comprehensive experimental framework to evaluate agent-to-agent negotiation and decision-making. (3) We conduct a large-scale analysis of several LLM-based agents and identify key risk factors that can lead to economic losses in autonomous real-world transactions.
@@ 2 Modeling Agent-to-Agent Negotiations and Transactions
The goal of this paper is to systematically investigate the outcomes and risks when AI agents are authorized to negotiate and make decisions on behalf of consumers and business owners. To this end, we introduce an experimental setting that closely reflects real-world negotiation and transaction scenarios in consumer markets. More specifically, we instruct LLM agents to engage in price negotiations over real consumer products, with one agent acting as the buyer and the other as the seller. By observing model behaviors in these structured and realistic scenarios, we aim to forecast potential behaviors, strategies, and risks that may arise as such agent-mediated transactions become more prevalent in future consumer environments.
@@ 2.2 Negotiation Scenario
In our negotiation simulation, buyer-seller interactions form an information-incomplete and zero-sum game (harsanyi1995games; raghavan1994zero; bianchi2024well). Both parties observe the item’s retail price $p_{r}$, but only the seller has access to the wholesale cost $p_{w}$. The buyer is permitted to accept, reject offers or continue to next round negotiation based on its budget $\beta$, while both agents are subject to strict feasibility constraints: No agreement may be reached if the final transaction price falls below the wholesale cost $p_{w}$ (for the seller) or exceeds the buyer’s budget $\beta$. We introduce the buyer’s budget $\beta$ to mirror real-world delegation scenarios, where users authorize buyer agents to act on their behalf within specified financial limits, such as account balances or spending caps. Within this setting, agents iteratively exchange offers and counteroffers to reach an agreement. The seller aims to keep the price close to retail, while the buyer attempts to maximize their discount.
@@ 2.3 Negotiation Pipeline
The negotiation is initiated by the buyer agent, who is required to open the conversation with an expression of interest in the product and a first offer (see greeting prompt for buyer in Appendix F.2). Then the two agents take turns to continue this negotiation until a termination condition is met. We use GPT-4o as the judge to decide whether a deal has been made by the buyer and the seller. At each round $t$, this judge model analyzes the buyer’s response and outputs a decision $d_{t}$, where $d_{t}\in\{\texttt{accept},\texttt{reject},\texttt{continue}\}$, indicating whether the buyer accepts the deal, rejects the negotiation entirely, or proceeds to the next round. The negotiation terminates immediately once $d_{t}$ is either accept or reject (see detailed prompt in Appendix F.4). To prevent excessively long interactions, we impose a maximum round limit of $T_{\max}$. Negotiations that reach this limit without resolution are treated as rejections, with the final decision $d_{T}$ set to reject. Moreover, if the final decision $d_{T}$ is accept, the proposed price in that round is recorded as the final transaction price.
@@ 2.4 Real-World Product Dataset
We construct a dataset $\mathcal{D}$ with 100 real consumer products drawn from three categories: motor vehicles, electronic devices, and real estate. To mimic real-world consumer settings, we collect the real retail price $p_{r}$ and key features $\phi$ for each item from trustworthy sources. As the wholesale cost $p_{w}$ may not be directly available on the public internet, we prompt GPT-4o with item-specific information and current market conditions to estimate a reasonable wholesale cost $p_{w}$ based on industry norms. More details of dataset creation are shown in Appendix D.
@@ 2.5 Agents Roles Design
To design agents that mimic real business negotiation settings, we construct the system prompts for each agent with the following four types of information: (1) Background: The background information of the agent. The seller is given $\{p_{r},p_{w},\phi\}$, while the buyer receives $\{p_{r},\beta,\phi\}$. (2) Goal: Both agents are asked to optimize the final price $p_{a}^{T}$ with respect to the retail price $p_{r}$. The seller seeks to maximize the profit, while the buyer is instructed to obtain the highest discount rate. (3) Constraint: The agents are instructed to follow certain constraints depending on their roles. For the seller agent, if the final decision $d_{T}$ is accept, the seller must comply with $p_{a}^{T}\geq p_{w}$, ensuring the final accepted price stays above the wholesale cost. The buyer is constrained by $p_{a}^{T}\leq\beta$ to follow budget limitations. Also, agents are instructed to reject a deal when facing an invalid agreement. (4) Guideline: A rule set governs interaction protocols that ensures agents follow realistic negotiation conventions. For example, buyers should avoid revealing their maximum budget in most situations, while sellers should avoid disclosing their wholesale price directly. Detailed system prompts of both agents can be found in Appendix F.1 & F.3.
@@ 2.6 Metrics
To quantify model negotiation performances, we created two primary metrics: (1) Price Reduction Rate ($PRR$), which measures a buyer model’s ability to negotiate discounts from the retail price $p_{r}$. Given the zero-sum nature of the game, $PRR$ also reflects seller performance, as a lower $PRR$ suggests greater success in resisting price reductions. (2) Relative Profit ($RP$), which directly measures a model’s capability to generate profit given a fixed set of products. Due to the large price difference among the three product categories, we present each model’s profit relative to the lowest-profit seller in the same setting. To further analyze sellers’ negotiation tendency, we also report two auxiliary metrics: Profit Rate (the average revenue per completed transaction) and Deal Rate (the proportion of negotiations that end successfully). These two metrics do not directly reflect an agent’s negotiation capability. Detailed mathematical formulas of metrics can be found in Appendix E.1.
@@ 3.1 Experimental Setup
We run our evaluations using nine LLM agents, including GPT series (o3, o4-mini, GPT4.1, GPT-4o-mini and GPT-3.5) (hurst2024gpt), DeepSeek series (DeepSeek-v3 (liu2024deepseek) and DeepSeek-R1 (guo2025deepseek)), and Qwen2.5 series (7B and 14B) (yang2024qwen2).To eliminate positional bias, we design the experiments with each model playing both the buyer and seller roles, interacting with every other model–including itself. To mimic real consumer settings, we define five discrete buyer budget levels, as shown in Table 1. These budget levels are intentionally varied to capture a wide spectrum of negotiation conditions–including under-constrained settings (where the buyer has ample budget), tightly constrained settings, and even economically irrational scenarios where the budget $\beta$ falls below the wholesale cost $p_{w}$. For evaluation, we randomly sample 50 products, and for each product, we run five trials, one per budget configuration. Furthermore, we set the maximum number of negotiation rounds, $T_{\text{max}}=30$.
3.2 Main Results
@@
Figure 2: Left: $PRR$ for both buyer and seller. Models located in the top-right region exhibit stronger relative negotiation performance, characterized by greater ability to push prices down when acting as buyers and to maintain higher prices when acting as sellers, reflecting overall bargaining power. Right: Seller agents’ relative profit rate, deal rate, and total profits.
@@
Disparity in Negotiation Capability Across Models. Given the zero-sum nature of our setting, the $PRR$ serves as a direct indicator of a model’s negotiation strength, capturing its performance both as a buyer and a seller. As illustrated in Figure 2 (Left), models exhibit substantial disparities in negotiation capabilities. Notably, o3 stands out with the strongest overall negotiation performance—demonstrating exceptional price retention as a seller and achieving the highest discount rate as a buyer. GPT-4.1 and o4-mini follow closely behind. In contrast, GPT-3.5 consistently underperforms across both roles, indicating the weakest negotiation ability among the models evaluated.
@@
The Trade-off Between Deal Rate and Profit Rate. To further assess models’ performance and behavior as seller agents, Figure 2 (Right) presents the seller-side metric—Relative Profit ($RP$)—the models’ total profit relative to the weakest model in our setting (i.e., GPT3.5), as well as the models’ average profit rate and deal rate. Most models outperform GPT-3.5 by approximately 9.6× in total profit, with GPT-4.1 and DeepSeek-R1 achieving 13.3× and 12×, respectively, leading all models. Notably, high-performing sellers such as o4-mini, GPT-4.1, and DeepSeek-R1 effectively balance profit margins with deal success rates, resulting in superior $RP$ scores. In contrast, other models struggle to manage this trade-off: GPT-4o-mini achieves the highest profit rate but suffers from low deal completion, while Qwen2.5-7B/14B and GPT-3.5 complete more deals but at the cost of thin profit margins—ultimately yielding lower total profits.
Figure 3: Average Price Reduction Rate ($PRR$) for each agent pair.
Figure 4: Average Deal Rate of seller agents over 5 budget settings.
Asymmetric Impacts of Agent Roles. As shown in Figure 4, the heatmap illustrates the $PRR$ across all pairwise combinations of buyer and seller agents. Our analysis reveals a clear asymmetry in agent roles: the choice of the seller model has a significantly larger impact on negotiation outcomes than the choice of the buyer model. For example, when we fix the seller as GPT-3.5 and vary the buyer agents, the difference between the highest and lowest $PRR$ is only 2.6%. In contrast, when we fix the buyer as GPT-3.5 and vary the seller agents, the $PRR$ gap reaches up to 14.9%. This asymmetry also explains the observation in Figure 2 (Left), where the average $PRR$ across different buyer agents shows relatively small variance: models’ capabilities have a larger impact on seller agents, but have a smaller impact on buyer agents.
Deal Rates in Different Budget Settings. Does the buyer’s budget affect the seller’s strategy and the deal rates? As shown in Figure 4, stronger models like GPT-4.1, o4-mini, and DeepSeek-R1 dynamically adapt to different budget scenarios and effectively adjust deal rates based on negotiation dynamics. Conversely, GPT-4o-mini and o3 consistently underperform with below-average deal rates across all budget levels. Low transaction volume undermines total revenue despite any profit margin advantages (as shown in Figure 2) (Right). GPT-3.5 and Qwen2.5-7b maintain above-average deal rates in all settings, indicating aggressive trading strategies that secure deals but yield lower profit rates.
Figure 5: Qwen models with more parameters obtain better deals as both sellers and buyers when they are negotiating with each other (Left) and DeepSeek-R1 (Right).

### v4

# The Automated but Risky Game: Modeling and Benchmarking Agent-to-Agent Negotiations and Transactions in Consumer Markets

@@ Abstract
AI agents are increasingly used in consumer applications for product search, negotiation, and transactions. We investigate a setting where both consumers and merchants authorize AI agents to automate negotiations and transactions. We address two questions: (1) Do different LLM agents exhibit varying performance when making deals for users? (2) What are the risks when using AI agents to fully automate negotiations in consumer settings? We design an experimental framework to evaluate AI agents’ capabilities in real-world negotiation scenarios, experimenting with various open-source and closed-source LLMs. Our analysis reveals that deal-making with LLM agents is an inherently imbalanced game. Furthermore, LLMs’ behavioral anomalies might lead to financial losses for both consumers and merchants through overspending or unreasonable deals. While automation can enhance efficiency, it poses significant risks to consumer markets. Users should be cautious when delegating business decisions to LLM agents. All the code and data are available at https://github.com/ShenzheZhu/A2A-NT.
The Automated but Risky Game: Modeling and Benchmarking Agent-to-Agent Negotiations and Transactions in Consumer Markets
Shenzhe Zhu1,2, Jiao Sun3, Yi Nian4, Tobin South5, Alex Pentland1,5, Jiaxin Pei1thanks: Correspondence to: cho.zhu@mail.utoronto.ca and pedropei@stanford.edu 1Stanford University  2University of Toronto  3Google DeepMind 4University of Southern California  5Massachusetts Institute of Technology
@@ 1 Introduction
Business negotiation and deal-making lie at the heart of the modern economy, yet achieving agreement is rarely straightforward. It requires effective information gathering, strategic reasoning, and skilled negotiation and decision-making (Lewicki and Hiam, 2011; Agndal et al., 2017). Recently, large language model (LLM) powered AI agents have demonstrated remarkable capabilities and are increasingly adopted for real-world tasks (Xu et al., 2024; Masterman et al., 2024). Given the importance of negotiation in business operations, researchers and practitioners have begun exploring ways to leverage AI agents to automate shopping and sales processes for both consumers and merchants (Kong et al., 2025; Chen et al., 2024), mostly assuming agents interact with real human users. However, with rapid AI agent adoption in consumer markets, both consumers and merchants might delegate their negotiation and decision-making to AI agents and direct agent-to-agent interactions might soon be commonplace. Given the natural capability differences of AI agents in negotiation settings (Bianchi et al., 2024) and unique agent-to-agent negotiation dynamics (Vaccaro et al., 2025), it becomes a key question: What happens when consumers and merchants use AI Agents with different capabilities to automate their negotiation and transactions in consumer settings?
@@
In this study, we propose a comprehensive framework to investigate opportunities and risks associated with fully automated, user-authorized agent-to-agent negotiation and transaction. Inspired by real-world shopping and sales workflows, we design an experimental setting where a buyer agent attempts to negotiate lower prices based on user-defined budgets, while a seller agent, aware of wholesale costs, aims to maximize profit. Each agent independently makes decisions throughout negotiation, simulating fully autonomous, end-to-end transactions between AI agents. To evaluate negotiation behaviors and capabilities of AI agents in realistic consumer scenarios, we compile a dataset of 100 real-world products across three major categories: electronic devices, motor vehicles, and real estate. These products vary in nature and price range, reflecting diverse consumer transactions. For each item, we collected actual retail prices and estimated wholesale values, which were provided to seller agents to simulate authentic market dynamics. We conducted negotiation and transaction experiments using several advanced language models, including GPT series (Hurst et al., 2024), Qwen-2.5 series (Yang et al., 2024a), and DeepSeek series (Liu et al., 2024; Guo et al., 2025). Our analysis reveals substantial negotiation performance gaps across models that correlate with their general capabilities and specifications. More capable models consistently secure better deals as both buyers and sellers. This suggests that in real-world scenarios, parties using less capable AI agents would face systematic economic disadvantages and financial losses.
@@
Beyond performance differences, we identify several key risks associated with delegating negotiation and transactional authority to AI agents: (1) Constraint violation risk: Buyer agents may disregard user-imposed budget constraints, completing purchases users cannot afford. Similarly, seller agents may accept prices below wholesale costs, leading to financial losses; (2) Excessive payment risk: buyer agents sometimes offer higher prices than retail price, resulting in unnecessary overpayment; (3) Negotiation deadlock risk: agents may become stuck in prolonged negotiation loops without reaching agreement; (4) Early settlement risk: higher budget settings lead buyer agents to compromise more readily, instead of striving for better deals. This contrasts with low-budget scenarios, where agents demonstrate stronger price resistance and negotiation effort.
@@
These findings have important implications for agent-assisted decision-making in consumer markets. Access to more powerful AI models can lead to better deals, potentially reinforcing economic disparities among users. Furthermore, weaknesses in LLMs, such as limited numerical reasoning and occasional failures in instruction-following, can expose both consumers and businesses to systemic financial risks. As fully autonomous agent-to-agent interactions become more common, practitioners should exercise caution when delegating high-stakes decisions to AI agents. This paper makes the following contributions:
  • We propose a novel and realistic setting for agent-to-agent negotiation and transaction, with clear practical implications for future consumer markets.
  • We design a comprehensive experimental framework to evaluate agent-to-agent negotiation and decision-making.
  • We conduct large-scale analysis of several LLM-based agents and identify key risk factors that can lead to economic losses in autonomous real-world transactions.
@@ 2 Modeling Agent-to-Agent Negotiations and Transactions
The goal of this paper is to systematically investigate outcomes and risks when AI agents are authorized to negotiate and make decisions on behalf of consumers and business owners. To this end, we introduce an experimental setting that closely reflects real-world negotiation and transaction scenarios in consumer markets. More specifically, we instruct LLM agents to engage in price negotiations over real consumer products, with one agent acting as buyer and the other as seller. By observing model behaviors in these structured and realistic scenarios, we aim to forecast potential behaviors, strategies, and risks that may arise as such agent-mediated transactions become more prevalent in future consumer environments.
@@ 2.2 Negotiation Scenario
In our negotiation simulation, buyer-seller interactions form an information-incomplete and zero-sum game (Harsanyi, 1995; Raghavan, 1994; Bianchi et al., 2024). Both parties observe the item’s retail price $p_{r}$, but only the seller has access to the wholesale cost $p_{w}$. The buyer is permitted to accept, reject offers or continue to next round negotiation based on its budget $\beta$, while both agents are subject to strict feasibility constraints: No agreement may be reached if the final transaction price falls below the wholesale cost $p_{w}$ (for the seller) or exceeds the buyer’s budget $\beta$. We introduce the buyer’s budget $\beta$ to mirror real-world delegation scenarios, where users authorize buyer agents to act on their behalf within specified financial limits, such as account balances or spending caps. Within this setting, agents iteratively exchange offers and counteroffers to reach an agreement. The seller aims to keep the price close to retail, while the buyer attempts to maximize their discount.
@@ 2.3 Negotiation Pipeline
The negotiation is initiated by the buyer agent, who is required to open the conversation with an expression of interest in the product and a first offer (see greeting prompt for buyer in Appendix E.2). Then the two agents take turns to continue this negotiation until a termination condition is met. In each round $t$, we deploy GPT-4o as an analyst to extract the most recent proposed price $p_{a}^{t}$ based on current round dialogue (see detailed prompt in Appendix E.5). Also, GPT-4o plays as a judge to decide whether a deal has been made by the buyer and the seller. At each round $t$, this judge model analyzes the buyer’s response and outputs a decision $d_{t}$, where $d_{t}\in\{\texttt{accept},\texttt{reject},\texttt{continue}\}$, indicating whether the buyer accepts the deal, rejects the negotiation entirely, or proceeds to the next round. The negotiation terminates immediately once $d_{t}$ is either accept or reject (see prompt in Appendix E.4). To prevent excessively long interactions, we impose a maximum round limit of $T_{\max}$. Negotiations that reach this limit without resolution are treated as rejections, with the final decision $d_{T}$ set to reject. Moreover, if the final decision $d_{T}$ is accept, the proposed price in that round is recorded as the final transaction price.
@@ 2.4 Real-World Product Dataset
We construct a dataset $\mathcal{D}$ with 100 real consumer products drawn from three categories: motor vehicles, electronic devices, and real estate. To mimic real-world consumer settings, we collect the real retail price $p_{r}$ and key features $\phi$ for each item from trustworthy sources. As the wholesale cost $p_{w}$ may not be directly available on the public internet, we prompt GPT-4o with item-specific information and current market conditions to estimate a reasonable wholesale cost $p_{w}$ based on industry norms. More details of dataset creation are shown in Appendix C.
@@ 2.5 Agents Roles Design
To design agents that mimic real business negotiation settings, we construct the system prompts for each agent with the following four types of information: (1) Background: The background information of the agent. The seller is given $\{p_{r},p_{w},\phi\}$, while the buyer receives $\{p_{r},\beta,\phi\}$. (2) Goal: Both agents are asked to optimize the final price $p_{a}^{T}$ with respect to the retail price $p_{r}$. The seller seeks to maximize the profit, while the buyer is instructed to obtain the highest discount rate. (3) Constraint: The agents are instructed to follow certain constraints depending on their roles. For the seller agent, if the final decision $d_{T}$ is accept, the seller must comply with $p_{a}^{T}\geq p_{w}$, ensuring the final accepted price stays above the wholesale cost. The buyer is constrained by $p_{a}^{T}\leq\beta$ to follow budget limitations. Also, agents are instructed to reject a deal when facing an invalid agreement. (4) Guideline: A rule set governs interaction protocols that ensures agents follow realistic negotiation conventions. For example, buyers should avoid revealing their maximum budget in most situations, while sellers should avoid disclosing their wholesale price directly. Detailed system prompts of both agents can be found in Appendix E.1 & E.3.
@@ 2.6 Metrics
To quantify model negotiation performances, we created two primary metrics: (1) Price Reduction Rate ($PRR$), which measures a buyer model’s ability to negotiate discounts from the retail price $p_{r}$. Given the zero-sum nature of the game, $PRR$ also reflects seller performance, as a lower $PRR$ suggests greater success in resisting price reductions. (2) Relative Profit ($RP$), which directly measures a model’s capability to generate profit given a fixed set of products. Due to the large price difference among the three product categories, we present each model’s profit relative to the lowest-profit seller in the same setting. To further analyze sellers’ negotiation tendency, we also report two auxiliary metrics: Profit Rate (the average revenue per completed transaction) and Deal Rate (the proportion of negotiations that end successfully). These two metrics do not directly reflect an agent’s negotiation capability. Detailed mathematical formulas of metrics can be found in Appendix D.1.
@@ 3.1 Experimental Setup
We evaluate agents across nine models, including GPT series(o3, o4-mini, GPT4.1, GPT-4o-mini and GPT-3.5) (Hurst et al., 2024), DeepSeek series(DeepSeek-v3 (Liu et al., 2024) and DeepSeek-R1 (Guo et al., 2025)), and Qwen2.5 series(7B and 14B) (Yang et al., 2024a), which constitute the core models used in our experiments. To eliminate positional bias, we design the experiments with each model playing both the buyer and seller roles, interacting with every other model–including itself. We define five discrete buyer budget levels, as shown in Table 1. These budget levels are intentionally varied to capture a wide spectrum of negotiation conditions–including under-constrained settings (where the buyer has ample budget), tightly constrained settings, and even economically irrational scenarios where the budget $\beta$ falls below the wholesale cost $p_{w}$. For evaluation, we randomly sample 50 products, and for each product, we run five trials, one per budget configuration. Furthermore, we set the maximum number of negotiation rounds, $T_{\text{max}}=30$.
3.2 Benchmark Results
@@
Figure 2: Top: PRRPRR for both buyer and seller. Models located in the top-right region exhibit stronger relative negotiation performance, characterized by greater ability to push prices down when acting as buyers and to maintain higher prices when acting as sellers, reflecting overall bargaining power. Bottom: Seller agents’ relative profit rate, deal rate, and total profits.
Disparity in Negotiation Capability Across Models.
@@
Given the zero-sum nature of our setting, $PRR$ serves as a direct indicator of a model’s negotiation strength, capturing its performance both as buyer and seller. As illustrated in Figure 2 (top), models exhibit substantial disparities in negotiation capabilities. Notably, o3 stands out with the strongest overall negotiation performance—demonstrating exceptional price retention as seller and achieving the highest discount rate as buyer. GPT-4.1 and o4-mini follow closely behind. In contrast, GPT-3.5 consistently underperforms across both roles, indicating the weakest negotiation ability among the models evaluated.
The Trade-off Between Deal Rate and Profit Rate.
@@
To further assess models’ performance and behavior as seller agents, Figure 2 (bottom) presents the seller-side metric—RPRP—which is computed relative to the total profit achieved by GPT-3.5, the model with the lowest absolute profit in our setting. Two additional indicators—average profit rate and deal rate—are also included to support the comparison. Most models outperform GPT-3.5 by approximately 9.6× in total profit, with GPT-4.1 and DeepSeek-R1 achieving 13.3× and 12×, respectively, leading all models. Notably, high-performing sellers such as o4-mini, GPT-4.1, and DeepSeek-R1 effectively balance profit margins with deal success rates, resulting in superior $RP$ scores. In contrast, other models struggle to manage this trade-off: GPT-4o-mini achieves the highest profit rate but suffers from low deal completion, while Qwen2.5-7B/14B and GPT-3.5 complete more deals but at the cost of thin profit margins—ultimately yielding lower total profits.
3.3 Agents’ Negotiation Capability Scales with Model Size